At what rate of simple interest will the sum double itself in 2 years?

Grade:8

2 Answers

ROSHAN MUJEEB
3 years ago
Dear Student
Simple interest formula is

A=P (1+rt)
Where p = principal or sum
r= rate of interest
t = time period
A = total return or principal plus interest earned during the time period.
So simply put your inputs in the equation I.e.
A = 2 X principal
Time frame = 2 years
We get,
2P = P (1+r2)
Or 2= 1+ r2 or r = (2–1)/2
r =50%
RISHIKA
3 years ago

 

HI FRIEND

The answer is 50% and the calculations are mentioned below :-

Simple interest formula is

A=P (1+rt)

Where p = principal or sum

r = rate of interest

t = time period

A = total return or principal plus interest earned during time period.

So simply putting your inputs in the equation I.e.

A = 2 X principal

Time frame = 2 years

We get,

2P = P (1+r2)

Canceling p as they are common, we get

Or 2= 1+ r2 or r = (2–1)/2

r =1/2 or 50%

Hope it helps!

 

THANKS